# LEMUR Neural Network Dataset: Towards Seamless AutoML

### v2

@@ 8 Determined Accuracies
Model Params. (×$\times 10^{6}$) Res. (px) Accuracy
AirNet 4.91 32$\times$32 0.8077
AirNext 1.51 256$\times$256 0.7769
AlexNet 57.04 299$\times$299 0.8675
BagNet 1.25 512$\times$512 0.6824
BayesianNet-1 3.57 32$\times$32 0.6710
BayesianNet-2 4.35 32$\times$32 0.6046
BayesianNet-3 0.12 32$\times$32 0.5509
ComplexNet 0.52 32$\times$32 0.7268
ConvNeXt 49.46 32$\times$32 0.5617
DPN107 0.02 32$\times$32 0.7340
DPN131 0.06 32$\times$32 0.7685
DPN68 0.06 32$\times$32 0.6530
DarkNet 3.66 32$\times$32 0.8499
DenseNet 25.53 128$\times$128 0.8792
EfficientNet 4.02 256$\times$256 0.9274
GoogLeNet 9.96 299$\times$299 0.9182
ICNet 0.05 256$\times$256 0.7166
InceptionV3-1 21.81 512$\times$512 0.8606
InceptionV3-2 24.37 512$\times$512 0.8665
MNASNet 1.9 299$\times$299 0.8280
MaxVit 30.38 299$\times$299 0.8813
MobileNetV2 2.24 299$\times$299 0.8661
MobileNetV3 0.59 512$\times$512 0.8688
RegNet 3.91 299$\times$299 0.8495
ResNet 11.18 299$\times$299 0.8370
ShuffleNet 1.26 299$\times$299 0.8441
SqueezeNet-1 0.74 299$\times$299 0.8005
SqueezeNet-2 0.73 32$\times$32 0.6913
SwinTransformer 27.53 32$\times$32 0.7482
UNet2D 0.56 32$\times$32 0.5796
VGG 128.81 299$\times$299 0.8242
VisionTransformer 85.23 32$\times$32 0.4885
@@
Model Params. (×$\times 10^{6}$) Res. (px) mIoU
DeepLabV3-1 39.84 64$\times$64 0.3731
DeepLabV3-2 58.83 128$\times$128 0.3543
FCN16s 15.31 256$\times$256 0.4365
FCN32s-1 32.95 512$\times$512 0.4386
FCN32s-2 51.94 128$\times$128 0.3200
FCN32s-3 15.31 512$\times$512 0.4101
FCN8s 15.31 64$\times$64 0.3020
LRASPP 3.22 256$\times$256 0.3810
UNet-1 31.04 128$\times$128 0.4494
UNet-2 17.26 64$\times$64 0.3393
@@
Model Params. (×$\times 10^{6}$) Res. (px) mAP
FCOS 32.13 800$\times$1333 0.7394
FasterRCNN 41.37 800$\times$1333 0.6214
RetinaNet 32.31 800$\times$1333 0.0922
SSDLite 3.17 300$\times$300 0.3975

### v4

@@ 8 Determined Accuracies
Model Params. ($\times 10^{6}$) Res. (px) Accuracy
AirNet 4.91 $\times32$ 0.8077
AirNext 1.51 $\times256$ 0.7769
AlexNet 57.04 $\times299$ 0.8675
BagNet 1.25 $\times512$ 0.6824
BayesianNet-1 3.57 $\times32$ 0.6710
BayesianNet-2 4.35 $\times32$ 0.6046
BayesianNet-3 0.12 $\times32$ 0.5509
ComplexNet 0.52 $\times32$ 0.7268
ConvNeXt 49.46 $\times32$ 0.5617
DPN107 0.02 $\times32$ 0.7340
DPN131 0.06 $\times32$ 0.7685
DPN68 0.06 $\times32$ 0.6530
DarkNet 3.66 $\times32$ 0.8499
DenseNet 25.53 $\times128$ 0.8792
EfficientNet 4.02 $\times256$ 0.9274
GoogLeNet 9.96 $\times299$ 0.9182
ICNet 0.05 $\times256$ 0.7166
InceptionV3-1 21.81 $\times512$ 0.8606
InceptionV3-2 24.37 $\times512$ 0.8665
MNASNet 1.9 $\times299$ 0.8280
MaxVit 30.38 $\times299$ 0.8813
MobileNetV2 2.24 $\times299$ 0.8661
MobileNetV3 0.59 $\times512$ 0.8688
RegNet 3.91 $\times299$ 0.8495
ResNet 11.18 $\times299$ 0.8370
ShuffleNet 1.26 $\times299$ 0.8441
SqueezeNet-1 0.74 $\times299$ 0.8005
SqueezeNet-2 0.73 $\times32$ 0.6913
SwinTransformer 27.53 $\times32$ 0.7482
UNet2D 0.56 $\times32$ 0.5796
VGG 128.81 $\times299$ 0.8242
VisionTransformer 85.23 $\times32$ 0.4885
@@
Model Params. ($\times 10^{6}$) Res. (px) mIoU
DeepLabV3-1 39.84 $\times64$ 0.3731
DeepLabV3-2 58.83 $\times128$ 0.3543
FCN16s 15.31 $\times256$ 0.4365
FCN32s-1 32.95 $\times512$ 0.4386
FCN32s-2 51.94 $\times128$ 0.3200
FCN32s-3 15.31 $\times512$ 0.4101
FCN8s 15.31 $\times64$ 0.3020
LRASPP 3.22 $\times256$ 0.3810
UNet-1 31.04 $\times128$ 0.4494
UNet-2 17.26 $\times64$ 0.3393
@@
Model Params. ($\times 10^{6}$) Res. (px) mAP
[…]
SSDLite 3.17 $\times300$ 0.3975